if it is given that f(x)=x^2-5x-14,please help me to determine the value of k, for which f(x)=k will have two unequal positive real roots
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

if it is given that f(x)=x^2-5x-14,please help me to determine the value of k, for which f(x)=k will have two unequal positive real roots
----------
If the discriminant is > 0, there are 2 real roots. Not necessarily positive.
---
f(x)=x^2-5x-14 = k
x^2-5x-14 - k = 0
-----
Find the discriminant.
===================
PS Helping you does not mean doing it for you.

Answer by stanbon(75887)

if it is given that f(x)=x^2-5x-14,please help me to determine the value of k, for which f(x)=k will have two unequal positive real roots
Solve::
x^2-5x-14 = k
x^2-5x-(14+k) = 0
----
Note:: b^2-4ac must be > 0
---------------------
25-4*1(-14-k) > 0
25+56+4k > 0
4k > -81
k > -81/4
-------
Cheers,
Stan H.
--------